OpenStudy (anonymous):

Solve the equation for X. -5x-13=-8x-4

OpenStudy (anonymous):

wouldnt it be x=-3?

OpenStudy (anonymous):

-5x-13=-8x+4 +13 +13 -5x=-8x+9 +8 +8 3x=9 /3 /3 x=3

OpenStudy (anonymous):

oh yeah a positive cause - divided by a - = a positive right?

OpenStudy (anonymous):

yeah